OpenStudy (anonymous):

Write the equation of a line with the given information in point-slope form, through point (-2,3) and slope -1/2

OpenStudy (anonymous):

Do you know the point-slope formula?

OpenStudy (anonymous):

Uh.. y-y1=m(x-x1)?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

So m is the slope, and (x1,y1) is a point on the line.

OpenStudy (anonymous):

In this case you have m = -1/2, and (x1,y1) = (-2,3)

OpenStudy (anonymous):

So plug in those values into your formula

OpenStudy (anonymous):

Im confused. What am i plugging in?

OpenStudy (anonymous):

The m and the (x1,y1) you were given.

OpenStudy (anonymous):

They said the slope (m) is -1/2, So put that in for m. They said a point (x1,y1) on the line was (-2,3) so plug in those numbers for x1 and y1.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yes. so now i have y-3=-1/2(x+2)?

OpenStudy (anonymous):

Yes indeed. except I'd like to see the parens around the fraction just to be totally clear: y-3=(-1/2)(x+2) or with the editor: \[y-3 =- \frac{1}{2}(x+2)\]

OpenStudy (anonymous):

Then if you need to have it in slope-intercept form you just solve for the y.

OpenStudy (anonymous):

Well.. i dont know what i need it in. What is the equation of a line? Did i already do that?

OpenStudy (anonymous):

This is _an_ equation for the line.

OpenStudy (anonymous):

There are a lot of other equations that make this same line.

OpenStudy (anonymous):

Well.. then i will just keep it as is, and if tis wrong, too bad. (:

OpenStudy (anonymous):

it's not wrong. It's definitely an equation for that line. It just might not be the equation the teacher is looking for. But in that case they should be more specific.